OpenStudy (anonymous):

Let R be a region delimited by x-axis, y=sqroot(x) and y=2-x. Rotate R around y-axis. Calculate volume. WHY IS MY SETUP WRONG? :S

OpenStudy (anonymous):

\[\int\limits_{0}^{1} \pi ((2-y)^2-y)dy\]

OpenStudy (phi):

\[ y = \sqrt{x} \] means \[ x= y^2 \] you want x to be the radius of a disc, with area \( \pi x^2 \) and volume \( \pi x^2 dy\) \[x^2 = y^4 \] so you want \( -\pi y^4 dy\) not just \( -\pi y \ dy \)

OpenStudy (anonymous):

that's all there is to it?

OpenStudy (phi):

I think so. It should be \[ \int\limits_{0}^{1} \pi ((2-y)^2-y^4) \ dy \]